UT 20150505 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson
Observer AO: DMiller (LBTO)
Telescope Operator: GBechetti
LBTO Support: TSargent

Plan for the night

Weather is presumably bad, so we will continue with closed-dome ACF and repro-tests

Summary

Bad weather all night

Repro- and ACF-tests

Details

All times are UT
Files are luci2.20150505.NNNN.fits

02:10 Sunset, cloudy, T = +1deg, Hum = 100%, strong wind (about 15-18 m/s), some snow showers

Problems with the CalUnit which must be moved in manually (by JMorris)
The reason is a not working interlock for the ARGOS dichroic, which will hopefully
be fixed tomorrow. For the moment beeing JMorris made a temporary fix.

05:00 Weather remains the same, cloudy, 100% humidity and wind now above limits (24m/s)

### Repro-tests

All grating tests use 0.5" longslit, so move it in first @ROT=210

Grating TILT reproducibility tests...
Grat_210_tilt_repro_180.xml

### G200 ###
0014 0019 Grat_200_tilt_repro_000.xml
0020 0025 Grat_200_tilt_repro_090.xml
0026 0031 Grat_200_tilt_repro_180.xml
0032 0037 Grat_200_tilt_repro_270.xml
Error tilting grating @1.93um, resubmitted script worked

### G210 ###
0038 0043 Grat_210_tilt_repro_270.xml
0044 0049 Grat_210_tilt_repro_180.xml
0050 0055 Grat_210_tilt_repro_090.xml
0056 0061 Grat_210_tilt_repro_000.xml

### G040 ###
0062 0067 Grat_40_tilt_repro_000.xml
Error tilting grating @2.17um, resubmitted script worked
Just right click and select "continue" for recoverable problems
0068 0073 Grat_40_tilt_repro_090.xml
0074 0079 Grat_40_tilt_repro_180.xml
0080 0085 Grat_40_tilt_repro_270.xml

Loading a new XML script takes: 40, 33, 41, 37, 37, 35 seconds
No "type ahead" possibility
Queue Actions interactions:
queue should be resettable so we start out empty each afternoon
hopefully the redraw will be faster when queue empty
queue listed should be from newest to oldest
OEP and Scriptor need to get useful information into queue

Grating SELECT reproducibility tests...
0086 0094 Grat_select_repro_270.xml
0095 0103 Grat_select_repro_180.xml
0104 0112 Grat_select_repro_090.xml
0113 0121 Grat_select_repro_000.xml

Reconfigure for APramskiy testing of AFC...

Weather bad for rest of night (clouds, humidity and snow)

-- %USERSIG{JochenHeidt - 2015-05-05}%
Topic revision: r5 - 19 May 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback